OpenStudy (anonymous):

HELP ! g(x) = arctanh (log x) Solve for g'(x) and find values for x for which g'(x) is defined .

OpenStudy (anonymous):

thats easy

OpenStudy (anonymous):

\[\frac{d}{dx} \tanh^{-1} (x) = \frac{1}{1-x^2}\]

OpenStudy (anonymous):

u= ln(x) du/dx= (1/x)

OpenStudy (anonymous):

by chain rules \[\frac{d}{dx} \tanh^{-1} (\ln(x) ) = \frac{1}{1-(\ln(x))^2} \times \frac{1}{x}\]

OpenStudy (anonymous):

so the derivative is not defined when the denominator is zero

OpenStudy (anonymous):

so when x=0 , or when 1-[ln(x)]^2 =0

OpenStudy (anonymous):

ln(x) = +- 1 so not defined when \[x=0, \frac{1}{e} , e \]

OpenStudy (anonymous):

thanks ! you're awesome ! :D